OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. -334 -274 -284 -346

OpenStudy (anonymous):

would the answer be -284?

OpenStudy (anonymous):

no :s

OpenStudy (campbell_st):

well you need to find the common difference -226 = 26 + (22 -1) d - 252 = 21d solve for d to find the common differrence. next use the formula for a term in an arithmetic series \[T_{31} = 26 + (31 -1)d\] and evaluate

OpenStudy (anonymous):

now i got -334........ is that correct

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

okay thank you